Add.
4/ s-2 + 4/ s+2
(Simplify the answer)

Answer by jorel1380(3719)

4/ s-2 + 4/ s+2=
4(s+2)/(s+2)(s-2)+4(s-2)/(s+2)(s-2)=
4s+8+4s-8/s2-4=8s/s2-4..